# Quantum education in the frontier city of Kharkiv

### 22 Feb 2024 v1

@@ Abstract
keywords:
School of Physics and Technology, quantum education challenges, quantum mechanics, Russian aggression, recent quantum initiatives and prospects
1 Introduction
@@
1.1 Brief history of quantum education at the School of Physics and Technology, Kharkiv University
@@ 1.2 The essence of education at SPT
1.3 The achievements of SPT graduates and students
@@
2 Education during the war
2.1 Consequences of the full-scale Russian aggression against Ukraine for SPT
The SPT campus is located separately from the main building of the Karazin University. It stands on the Northern outskirt of Kharkiv city, in other words, it is one of the closest city structures to the Ukrainian state boarder. The building was partly destroyed by Russian missiles already during the first night of the Russian invasion, on February 24-25, 2022. Not only research labs and lecture halls were destroyed, as shown on Fig. 1. Student dormitory was hit by Russian missiles as well. Some of SPT professors, engineers and students voluntary joined Ukrainian Armed Forces to defend the Motherland. About 60% of teaching staff of Karazin University had to move from Kharkiv to Western cities of Ukraine or abroad. But those who stayed in the city had to overcome this terrifying circumstances.
Figure 1: The state of the SPT educational building as of February 2024: the lecture hall, the view from East, the view from North, and the laboratory of molecular spectroscopy.
@@ 3 Quantum courses at SPT
It is worth noting that there are modern textbooks and monographs for several courses taught at SPT, which are authored by the School’s professors. For instance, the foundational textbook on Quantum Mechanics [2], as well as monographs related to the courses on Quantum Information Theory [3] and Quantum Statistical Mechanics [4].
4 Recently emerged initiatives and the future
@@
After resounding success of the quantum forum, we realized the necessity to keep high level of scientific exchange with foreign colleagues. By joining the efforts from three Kharkiv institutions, we organized the Kharkiv Quantum Seminar series, which are held on a regular basis in the virtual format since October 2023 [8]. The seminars are given dominantly by the speakers on the middle stages of their research careers, to a large extent, by the exceptionally motivated representatives of Ukrainian scientific diaspora (Denys Bondar, Sergey Frolov, Kyrylo Snizhko et al.). At the same time, there are also invited speakers, the world-leading specialists in quantum physics, such as Michael Berry, John Perdew, Franko Nori et al., who expressed their sincere interest in giving lectures in these seminars. In addition, there are weekly seminars at SPT, with speakers being mostly researchers on early stages of their scientific careers (dominantly formed by recent graduates from SPT) sharing their knowledge with colleagues and researchers in Kharkiv.
SPT has recently introduced new BSc educational program on “Cyberphysical nuclear technologies”. It contains classical quantum courses as well as newly introduced ones, such as “Quantum computer basics” and “Quantum algorithms”. Currently, we are on the stage of development of the Master-level educational program oriented on quantum computing and quantum technologies. In particular, by joint efforts of representatives of Taras Shevchenko Kyiv National University (Maksym Teslyk) and University of South Carolina (Yaroslaw Bazaliy) the SPT students are attending the new course on quantum information theory. Also, SPT students participated in the winter school on quantum computing “QHack UA Bootcamp 2024” [9], where their team won the first prize.
As a bitter pill for the quantum community in Ukraine and, especially, in the frontier city of Kharkiv, we were heavily affected to learn about the specific details of the resolution for the 2023 General Conference of the United Nations Educational, Scientific and Cultural Organization (UNESCO) and the 2023 General Assembly of the United Nations to proclaim 2025 the International Year of Quantum Science and Technology, in particular, under the tab “our partners” [10]. As of February 2024, the corresponding webpage indicates that their partners are, in particular, the “Russian Academy of Sciences” (RAS) and “Russian Physical Society” (RPS). Also, the academic community should be aware of the fact that about 200 rectors (presidents) of Russian universities, led by Professor Viktor Sadovnichy, Rector of Moscow State University, signed a letter that justifies, approves and encourages war in Ukraine [11].
@@ 5 Conclusion
[…]
Acknowledgements The authors deeply appreciate the Ukrainian Armed Forces for the possibility to continue teaching students and carrying out research. Comprehensive support of the democratic world to Ukraine in its struggle against Russian invasion is highly appreciated.

### 21 Mar 2024 v2

@@ Abstract
I Introduction
@@
I.1 Brief history of quantum education at the School of Physics and Technology, Kharkiv University
@@ I.2 The essence of education at SPT
I.3 The achievements of SPT graduates and students
@@
II Education during the war
II.1 Consequences of the full-scale Russian aggression against Ukraine for SPT
The SPT campus is located separately from the main building of the Karazin University. It stands on the Northern outskirt of Kharkiv city, in other words, it is one of the closest city structures to the Ukrainian state boarder. The building was partly destroyed by Russian missiles already during the first night of the Russian invasion, on February 24-25, 2022. Not only research labs and lecture halls were destroyed, as shown in Fig. 1. Student dormitory was hit by Russian missiles as well. Some of SPT professors, engineers and students voluntary joined Ukrainian Armed Forces to defend the Motherland. About 60% of teaching staff of Karazin University had to move from Kharkiv to Western cities of Ukraine or abroad. But those who stayed in the city had to overcome this terrifying circumstances.
Figure 1: The state of the SPT educational building as of March 2024: the lecture hall, the view from East, the view from North, and the laboratory of molecular spectroscopy.
@@ III Quantum courses at SPT
It is worth noting that there are modern textbooks and monographs for several courses taught at SPT, which are authored by the School’s professors. For instance, the foundational textbook on Quantum Mechanics [2], as well as monographs related to the courses on Quantum Information Theory [3], Quantum Statistical Mechanics [4], Quantum Scattering Theory [5] and Theory of Nuclear Reactions [6].
IV Recently emerged initiatives and the future
@@
After resounding success of the quantum forum, we realized the necessity to keep high level of scientific exchange with foreign colleagues. By joining the efforts from three Kharkiv institutions, we organized the Kharkiv Quantum Seminar series, which are held on a regular basis in the virtual format since October 2023 [10]. The seminars are given dominantly by the speakers on the middle stages of their research careers, to a large extent, by the exceptionally motivated representatives of Ukrainian scientific diaspora (Denys Bondar, Sergey Frolov, Kyrylo Snizhko, Artem Volosniev, Olexandr Isayev, Denis Seletskiy, et al.). At the same time, there are also invited speakers, the world-leading specialists in quantum physics, such as Steven M. Anlage, Tomasz Dietl, Gediminas Juzeliūnas, Michael Berry, Jan Kuneš, John Perdew, Franco Nori, Roger Melko, et al., who expressed their sincere interest in giving lectures in these seminars. In addition, there are weekly seminars at SPT, with speakers being mostly researchers on early stages of their scientific careers (dominantly formed by recent graduates from SPT) sharing their knowledge with colleagues and researchers in Kharkiv.
SPT has recently introduced new BSc educational program on “Cyberphysical nuclear technologies”. It contains classical quantum courses as well as newly introduced ones, such as “Quantum computer basics” and “Quantum algorithms”. Currently, we are on the stage of development of the Master-level educational program oriented on quantum computing and quantum technologies. In particular, by joint efforts of representatives of Taras Shevchenko Kyiv National University (Maksym Teslyk) and University of South Carolina (Yaroslaw Bazaliy) the SPT students are attending the new course on quantum information theory. Also, SPT students participated in the winter school on quantum computing “QHack UA Bootcamp 2024” [11], where their team won the first prize. After that, they participated as the “Kharkiv_team” in the international QHack 2024 event, where they won the first prize in the challenge “A Matter of Taste” and were among top-5 teams in the open hackaton competition [12].
As a bitter pill for the quantum community in Ukraine and, especially, in the frontier city of Kharkiv, we were heavily affected to learn about the specific details of the resolution for the 2023 General Conference of the United Nations Educational, Scientific and Cultural Organization (UNESCO) and the 2023 General Assembly of the United Nations to proclaim 2025 the International Year of Quantum Science and Technology, in particular, under the tab “our partners” [13]. As of March 2024, the corresponding webpage indicates that their partners are, in particular, the “Russian Academy of Sciences” (RAS) and “Russian Physical Society” (RPS). Also, the academic community should be aware of the fact that about 200 rectors (presidents) of Russian universities, led by Professor Viktor Sadovnichy, Rector of Moscow State University, signed a letter that justifies, approves and encourages war in Ukraine [14].
@@ V Conclusion
Acknowledgements.
The authors deeply appreciate the Ukrainian Armed Forces for the possibility to continue teaching students and carrying out research. Comprehensive support of the democratic world to Ukraine in its struggle against Russian invasion is greatly acknowledged.